Π polarizability Meeting Mar 23, 2022
09:30 EST (JLab time) Weekly Meeting
REMOTE during Covid-19
- 1 Connecting
- 2 Action Items
- 3 Preliminary Agenda
- 4 Minutes
- 4.1 Cables
- 4.2 Run Start Date
- 4.3 TAC Status
- 4.4 Amount of Empty Target Data to Take
- 4.5 Diamond Simulations
- 4.6 Software/Calibration
- 4.7 Scintillators
- 4.8 Translation Tables for Scalers
- 4.9 Low Voltage Feedback for the Chambers
- 4.10 Lead Shielding Alignment
- 4.11 CPP Trigger Tests
- 4.12 DAQ and Electronic Setup
- 4.13 Tagger Hodoscope Location 7
- 4.14 CTOF Trigger
- 4.15 Schedule for Chamber Data
Join ZoomGov Meeting
Meeting ID: 160 862 8791
Elton Smith is inviting you to a scheduled ZoomGov meeting. Topic: CPP Weekly Meeting Time: This is a recurring meeting Meet anytime Join ZoomGov Meeting https://jlab-org.zoomgov.com/j/1608628791?pwd=bXhtZjdNRi8xeXZqWFh5T1R3TDBFdz09 Meeting ID: 160 862 8791 Passcode: CPP One tap mobile +16692545252,,1608628791# US (San Jose) +16468287666,,1608628791# US (New York) Dial by your location +1 669 254 5252 US (San Jose) +1 646 828 7666 US (New York) +1 669 216 1590 US (San Jose) +1 551 285 1373 US 833 568 8864 US Toll-free Meeting ID: 160 862 8791 Find your local number: https://jlab-org.zoomgov.com/u/acZK7hDWFf Join by SIP firstname.lastname@example.org Join by H.323 18.104.22.168 (US West) 22.214.171.124 (US East) Meeting ID: 160 862 8791 Passcode: 112160
- Check with Sasha on steps for "replay"
- Work with Justin to create the appropriate translation tables for the scalers (Beni)
- Check with Chris about sensor voltage for the LV supplies (Beni)
- Develop a draft plan for Pb alignment (Ilya)
- Previous CPP meeting: Mar 16, 2022
- Previous CPP Software meeting: Π polarizability Software Meeting Mar 17, 2022
- Action Items
- CPP and NPP Run Conditions
- TAC Runs / flux normalization: TAC Operation Test
- Diamond simulations (Richard)
- MWPC Testing
- CPP and NPP Experiment Configuration
- CPP-NPP trigger
- DAQ and Electronic Setup (Sergey and Beni)
- Other business
- Next meeting
Present: Alex Austregesilo, Albert Fabrizi, Sergey Furletov, David Hornidge, Mark Ito (chair), Igal Jaegle, Nikhil Kalra, Ilya Larin, Rory Miskimen, Simon Taylor, Beni Zihlmann
There is a recording of this meeting (Passcode: n.1a#yZ9) on the ZoomGov website.
Beni reported that the LV and HV cables for the chambers were being installed yesterday. The signal cables are still in productions. He reminded us that we have 6 signal cables in hand that can be used to test one chamber at a time.
[Added in press, from Beni: the crimp tool loan mentioned at the last meeting was approved yesterday and has been shipped to the company. Also Chris Stanislav is connecting the LV boxes to the wire chambers as of this afternoon.]
Run Start Date
Beni reminded us that the run is scheduled to start June 8. We should not now that a budget has passed, we should not assume that there will be a CR-related delay in the start of the run.
Beni had a look at the lead glass TAC in the Hall. He led us through his logbook entry. All indications are that the device looks good. He mentioned that that TAC was last used with GlueX in November/December 2019.
Beni also mentioned Sascha Somov's plan to use a new, 3×3 lead tungstate array as a TAC for CPP and PrimEx. We had some discussion of the relative merits of lead glass TAC vs. lead tungstate TAC.
- possible leakage of showers from the sides of PbWO4 vs. large size of Pb-glass
- better intrinsic energy resolution of PbWO4
- possible need for calibration, block-to-block of a PbWO4 array
- TAC stage only has horizontal motion. Snake scan requires two-dimensional motion
- PbWO4 is a new detector
Rory suggested that Mark ask Sascha to come to tomorrow's CPP Design meeting to present his plan.
Amount of Empty Target Data to Take
- Simon has moved the kinematic fit of pairs from a plugin to a factory in the FMWPC library. It has not been tested on electron-positron pairs yet. Rory remarked that its location in this library is motivated by having results of the fits available for online e/π separation monitoring. Without the kinematic fit, this would be difficult due to the relatively poor momentum resolution for forward-going tracks.
- Simon is working on putting the survey numbers in. He has a version in a private HDDS directory.
The CTOF paddles are being installed in the Hall this week. Bobby Bunton 3D-printed mounting brackets, at Beni's request, so that mechanical support attachments are only made on the light guides and not to the scintillators or the phototube/base.
Translation Tables for Scalers
Beni has a meeting with Justin scheduled for Friday to discuss how this is to be done.
Low Voltage Feedback for the Chambers
Beni confirmed with Chris that the feedback system for regulating low voltage at the distribution box is hooked up.
Lead Shielding Alignment
The lead shielding in front of the first MWPC will be positioned, at the 1 mm level, from a study using rates, with beam on, of the MWPCs at several positions of the shield. Position adjustment is done by hand in the Hall, so Rapid Access will be necessary to carry out the program. Ilya will draft a run plan for this activity.
CPP Trigger Tests
Sergey reported that the CDC AI tests are done. Sascha is aware of this and he will start back on trigger tests. Ben Raydo may have to be pulled in at some point. Rory emphasized that this is on the critical path.
DAQ and Electronic Setup
Sergey reported that the CDC crate with the FADC-125s is mounted in its new location for use with the MWPCs. A new VME computer has been put in the crate. The crate has no power and it needs a optical fiber for read-out still.
[Added in press, from Beni: the VXS crate has now power. There is also AC 120V regular power available in the rack.]
Tagger Hodoscope Location 7
Mark mentioned that at yesterday's Hall D Staff meeting Tim Whitlatch mentioned that he was waiting on a decision from the CPP group on installing a plate, at location 7, for mounting tagger hodoscope counters. We think we need this. We will try to clarify the issue at tomorrow's Design Meeting.
[Added in press: Mark asked and Sascha agreed to attend the meeting.]
Rory asked about readiness of the various components needed to form a trigger from the downstream scintillators. Beni reported that all the components were in-hand or easily acquired.
- The PMTs have two outputs, one will to the FADCs, the other to a discriminator
- The discriminator has two outputs. One will go to a CAEN TDC, the other to the trigger logic.
- Trigger logic is housed in a NIM crate.
- The output of the trigger logic modules is sent over fiber to the GTP.
Schedule for Chamber Data
Alex asked about the schedule for reading out the chambers so that monitoring and reconstruction can start to be tested. Beni thought that it would be at least couple of weeks. Not only does the chamber read-out chain need to be established, but the CTOF trigger has to be working.